Give the closed-form solution of the recurrence relation where indicated using algebraic unrolling and answer the questions regarding the sequence that is denoted by the relation

Solve T(n) = n + 1 + T(n − 1), n ≥ 2, with the initial condition T(2) = 3
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

T(n)=T(n-1)+(n+1)

So,

T(n)=T(n-2)+(n+1)+(n)

So,

T(n)=T(n-3)+(n+1)+n+(n-1)

..

..

T(n)=T(2)+(n+1)+n+(n-1)+.....4

So,

T(n)=3+4+5+.....(n)+(n+1)

So,

T(n)=(n+1)*(n+2)/2-3

Kindly revert for any queries

Thanks.